\begin{thebibliography}{99}

\bibitem{Mardoyan:2006qd}
  L.~Mardoyan, A.~Nersessian, H.~Sarkisyan and V.~Yeghikyan,
  ``Dipole transitions and Stark effect in the charge-dyon system,''
  arXiv:cond-mat/0609768.
  %%CITATION = COND-MAT 0609768;%%
\bibitem{Tomilchik:1963}
  L.~M.~Tomilchik,
   Sov. JETP  {\bf 17} (1963) 111.
\bibitem{Strazhev:1975}
  V.~I.~Strazhev and L.~M.~Tomilchik: ``Electrodynamics with a Magnetic Charge'' (Nauka i Tekhnika Publ., Minsk 1975, in Russian).
  %\cite{Tolkachev:1983st}
\bibitem{Tolkachev:1983st}
  E.~A.~Tolkachev, L.~M.~Tomilchik and Y.~M.~Shnir,
   ``Modification Of Selection Rules As P Violating Effect In The Charge
  Monopole System. (In Russian),''
  Yad.\ Fiz.\  {\bf 38} (1983) 541.
  %%CITATION = YAFIA,38,541;%%
  
  %\cite{Tolkachev:1988mn}
\bibitem{Tolkachev:1988mn}
  E.~A.~Tolkachev, L.~M.~Tomilchik and Y.~M.~Shnir,
  ``P Parity Violation in the Dyogen Spectrum,''
  J.\ Phys.\ G {\bf 14} (1988) 1.
  %%CITATION = JPHGB,G14,1;%%
  %\cite{Shnir:1992iz}
  %\cite{Tolkachev:1989cr}
\bibitem{Tolkachev:1989cr}
  E.~A.~Tolkachev, L.~M.~Tomilchik and Y.~M.~Shnir,
  ``The Dyogen atom in an external field and parity nonconservation,''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 50} (1989) 275
  [Yad.\ Fiz.\  {\bf 50} (1989) 442].
  %%CITATION = SJNCA,50,275;%%
  %\cite{Tolkachev:1990rw}
\bibitem{Tolkachev:1990rw}
  E.~A.~Tolkachev, L.~M.~Tomilchik and Y.~M.~Shnir,
  ``Monopole-induced excitation of an atom,''
  Sov.\ J.\ Nucl.\ Phys.\  {\bf 52} (1990) 916
  [Yad.\ Fiz.\  {\bf 52} (1990) 1447].
  %%CITATION = SJNCA,52,916;%%
\bibitem{Shnir:1992iz}
  Y.~M.~Shnir, E.~A.~Tolkachev and L.~M.~Tomilchik,
   ``P violating magnetic monopole influence on the behavior of the atom like
  system in external fields,''
  Int.\ J.\ Mod.\ Phys.\ A {\bf 7} (1992) 3747.
  %%CITATION = IMPAE,A7,3747;%%
  %\cite{Goddard:1977da}
\bibitem{Goddard:1977da}
  P.~Goddard and D.~I.~Olive,
  ``New Developments In The Theory Of Magnetic Monopoles,''
  Rept.\ Prog.\ Phys.\  {\bf 41} (1978) 1357.
  %%CITATION = RPPHA,41,1357;%%
%\cite{Zhang:1988ab}
\bibitem{Zhang:1988ab}
  J.~z.~Zhang and Y.~c.~Qi,
   ``Wave Functions of Bound States of a Fermion and a Dirac Dyon and Matrix Elements in an External Electromagnertic Field''
  J.\ Math.\ Phys.\  {\bf 31} (1990) 1796.
  %%CITATION = JMAPA,31,1796;%%
  %\cite{Zhang:1990fk}
\bibitem{Zhang:1990fk}
  J.~z.~Zhang,
   ``Dipole Moments Of The Bound System Of A Charged Fermion And An Abelian
  Dyon,''
  J.\ Phys.\ G {\bf 16} (1990) 195.
  %%CITATION = JPHGB,G16,195;%%
  %\cite{Zhang:2002rk}
\bibitem{Zhang:2002rk}
  J.~z.~Zhang,
  ``Testing Abelian dyon - fermion bound system,''
  Phys.\ Lett.\ B {\bf 532} (2002) 215
  [arXiv:hep-th/0205124].
  %%CITATION = HEP-TH 0205124;%%
  %\cite{Bose:1985yr}
\bibitem{Bose:1985yr}
  S.~K.~Bose,
  ``Bound States Of A Charged Particle And A Dyon,''
  J.\ Phys.\ A {\bf 18} (1985) 1289.
  %%CITATION = JPAGB,A18,1289;%%
  %\cite{Bose:1986mc}
\bibitem{Bose:1986mc}
  S.~K.~Bose,
  ``Dyon Electron Bound States,''
  J.\ Phys.\ G {\bf 12} (1986) 1135.
  %%CITATION = JPHGB,G12,1135;%%
  %\cite{Kazama:1976fm}
\bibitem{Kazama:1976fm}
  Y.~Kazama, C.~N.~Yang and A.~S.~Goldhaber,
   ``Scattering Of A Dirac Particle With Charge Ze By A Fixed Magnetic
  Monopole,''
  Phys.\ Rev.\ D {\bf 15} (1977) 2287.
  %%CITATION = PHRVA,D15,2287;%%
%\cite{Kazama:1977um}
\bibitem{Kazama:1977um}
  Y.~Kazama,
   ``Electric And Magnetic Dipole Moments Of The Bound System Of A Dirac
  Particle And A Fixed Magnetic Monopole,''
  Phys.\ Rev.\ D {\bf 16} (1977) 3078.
  %%CITATION = PHRVA,D16,3078;%%

\end{thebibliography}

\documentclass[preprintnumbers,amsmath,amssymb]{revtex4}
\begin{document}
\title{Comment on ``Dipole transitions and Stark effect in the charge-dyon system,''\\ by L. Mardoyan et al. \cite{Mardoyan:2006qd}}
\author{E. A. Tolkachev and L. M. Tomilchik}
\altaffiliation{email address: tolkachev, tomilchik@dragon.bas-net.by}
\affiliation{Institute of Physics, 68 Nezavisimosti Avenue, Minsk 220072, Belarus}
\date{Oct. 4, 2006}
\begin{abstract}
\end{abstract}
\maketitle
In a recent e-print ({\tt cond-mat/0609768}, ``Dipole transitions and Stark effect in the charge-dyon system'', by L.~Mardoyan, A.~Nersessian, H.~Sarkisyan, V.~Yeghikyan, \cite{Mardoyan:2006qd}) the authors claim that "{\it the presence of monopole, besides providing the system with degenerated (on azimuth quantum number) ground state, makes possible the dipole transitions         which obey selection rules $(l'=l; m'=m, m'=m\pm 1)$, where $l,m$ are          respectively, orbital and azimuth quantum numbers}".
Also they calculated the linear and quadratic Stark effect in such a system and pointed out that "{\it the linear Stark effect in the ground state is proportional to azimuth quantum number, and to the sign of monopole number}" and "{\it the quadratic Stark effect in the ground state is independent on the signs of azimuth and monopole numbers}".

It is well known \cite{Tolkachev:1983st,Tolkachev:1988mn}
that modification of the selection rules in a
theory with monopole arises as a straightforward consequence of the
non-invariance of model under spatial reflection (P) which is a by-product of consistent consideration of magnetic monopoles, both Abelian and non-Abelian
\cite{Tomilchik:1963,Strazhev:1975,Goddard:1977da}.

It was also demonstrated on a simple model example of a bound non-relativistic system electron-dyon, which is described by Schr\"odinger or Pauli Hamiltonian, that the selection rules for a dipole radiation, unlike the familiar  quantum-mechanical situation, admit the transitions with $\Delta l =0$ \cite{Tolkachev:1988mn}. It was pointed out that in the consistent relativistic theory based on Dirac equation the same modification is taking place as well \cite{Kazama:1976fm,Kazama:1977um,Zhang:1990fk,Zhang:1988ab}.

Consideration of properties of the bound system charge-monopole or charge-dyon in external fields also allows to observe the above mentioned effects, which are   related with P-parity violation. In particular, Stark and Zeeman effects, as well as photoionization of the ground state of the  bound system charge-dyon was analyzed in \cite{Shnir:1992iz,Tolkachev:1989cr,Tolkachev:1990rw}. Since the modification of the selection rules for a dipole radiation arises from angular part of the Hamiltonian and is related with  matrix element sandwiched between generalized angular harmonics,  inclusion of additional radial-dependent interaction terms, does not affect this result. It particular, the spectrum of Pauli and Dirac Hamiltonian of such a system with usual Coulomb electromagnetic potential and corresponding matrix elements of the dipole moment operator were considered by \cite{Bose:1985yr,Bose:1986mc,Zhang:2002rk}.

The inclusion of the extra radial dependent term $s^2/2 r^2$ into the non-relativistic Hamiltonian of the charge-dyon system does not make any principal difference comparing to the previous consideration. At least, the authors have to refer the related works properly.

\end{document}